A random process has sample functions of the form X(t) = Acos(wt+Θ) where w is constant, A is a random variable that has a magnitude of +1 and −1 with equal probability, and Θ is a random variable that is uniformly distributed between 0 and 2π. Assume that the random variables A and Θ are independent. 1. Is X(t) a wide-sense stationary process?


1
Expert's answer
2020-08-31T17:09:24-0400

1.

"E[X(t)]=E[Acos(\omega t+\Theta)]."

Since "A" and "\Theta" are independent, then

"E[X(t)]=E[Acos(\omega t+\Theta)]=\\
E[A]\cdot E[cos(\omega t+\Theta)]=\\
(1\cdot1/2+(-1)\cdot1/2)E[cos(\omega t+\Theta)]=\\
0\cdot E[cos(\omega t+\Theta)]=0."

Expectation of "A^2":

"E[A^2]=(1^2\cdot1/2+(-1)^2\cdot1/2)=1."

2.

Correlation function:

"R(t_1,t_2)=E[X(t_1)X(t_2)]=\\
E[Acos(\omega t_1+\Theta)\cdot Acos(\omega t_2+\Theta)]=\\
E[A^2]E[cos(\omega t_1+\Theta)\cdot cos(\omega t_2+\Theta)]=\\"

"E[cos(\omega t_1+\Theta)\cdot cos(\omega t_2+\Theta)]=\\
E[\frac{1}{2}cos(\omega t_1+\Theta+\omega t_2+\Theta)+\\
+\frac{1}{2}cos(\omega t_1+\Theta-\omega t_2-\Theta)]=\\
E[\frac{1}{2}cos(\omega (t_1+ t_2)+2\Theta)+\frac{1}{2}cos(\omega( t_1- t_2))]=\\
E[\frac{1}{2}cos(\omega (t_1+ t_2)+2\Theta)]+\\
+E[\frac{1}{2}cos(\omega( t_1- t_2))]."

Since "sin(x)" is periodic, then

"E[\frac{1}{2}cos(\omega (t_1+ t_2)+2\Theta)]=\\
\intop_{0}^{2\pi}\frac{1}{2}cos(\omega (t_1+ t_2)+2\Theta)\frac{1}{2\pi}d\theta=\\
\frac{1}{4\pi}\frac{sin(\omega (t_1+ t_2)+2\Theta)}{2}|_0^{2\pi}=\\
\frac{1}{4\pi}\frac{sin(\omega (t_1+ t_2)+4\pi)}{2}-\frac{1}{4\pi}\frac{sin(\omega (t_1+ t_2))}{2}=0,"

hence

"R(t_1,t_2)=E[X(t_1)X(t_2)]=\\
E[\frac{1}{2}cos(\omega (t_1+ t_2)+2\Theta)]+\\
+E[\frac{1}{2}cos(\omega( t_1- t_2))]=\\
0+\frac{1}{2}cos(\omega( t_1- t_2))=\\
\frac{1}{2}cos(\omega( t_1- t_2))"

3.

Since "E[X(t)]=0" and "R(t_1,t_2)=\frac{1}{2}cos(\omega( t_1- t_2))" is not random and depends only on "(t_1-t_2)", then "X(t)" is a wide-sense stationary process.

Answer: Yes, "X(t)" is a wide-sense stationary process.